Substituting x = – 3 in
x^2 + 5x gives (– 3)^2 +5(–3) = – 9–15 = – 24

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Substituting x =–3 in x^2+5x
x^2+5x = (–3)^2+5(–3) = 9–15 = -6.
This is the incorrect answer

Thanks